A sample of lawn fertiliser was analysed by back titration to determine its ammonium ion content. The sample was boiled in NaOH solution. The reaction occurred according to the equation
NH4+(aq) + NaOH(aq)
1
Expert's answer
2020-08-03T07:41:10-0400

 The reaction between acids and aqueous ammonia forms ammonium ion


NH(aq)H+ (aq) = NH > 4 + (aq)

The ammonium ion and alkali metals ion behaves chemically the same particularly potassium ions .chemically like the ions of the Also all ammonium salts are white andalkali metals, particularly soluble.


Characteristic reactions of NH₄⁺Sodium Hydroxide


Addition of concentrated hydroxide ion solutions evolves NH3 gas:

NH4(aq) + OH(aq) = NH(aq) + H2O (l)

The ammonia gas turns moistened red litmus paper blue. It can also be detected by its odor. This reaction serves as a confirmatory test for NH4+